Berts age plus twice Ernies age is 31. Three times Berts age plus five times Ernies age is 85. How old is Bert and Ernie?
Answer by Cromlix(4381)

Bert's age = x
Ernie's age = y
x + 2y = 31.......1
3x + 5y = 85......2
Multiply Eq.(1) by 3
3x + 6y = 93......3
3x + 5y = 85......2
Subtract (2) from (3)
y = 8
Substitute y = 8 into
Eq(1).
x + 2(8) = 31
x + 16 = 31
x = 31 - 16
x = 15
Bert is 15 years old.
Ernie is 8 years old.
Hope this helps.
:-)